CH/Agen/94/01

Agenda

Charissa Group Meeting

To be held on: Wednesday, 5 Jan 1994, 1.30 pm

To be held at: York Room PT003 New Concourse Area

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

Agenda

MEGHA Meeting, 11 a.m., Room PT003, New Concourse Area Buffet Lunch 12.20 p.m.

1. Minutes of 17 November 1993
2. Matters Arising / Actions
3. Update on NPC/SERC matters  (DLW)
4. Progress with MEGHA development

Module bodies and feedthroughs

Scintillators

Grids and windows

Silicon mounting and purchasing

5. Progress with electronics construction  (WDMR,DLW,RAH)

Preamplifiers Amplifiers Power Supplies

6. Progress with Data Acquisition for next ANU running  (RAH,VP,SJAH)
7. Experiments and tests planned for next ANU running
8. Status of equipment for next ANU running

Refer: Ch/Doc/93/14, “Responsibilities for March 1994”

9. Scheduling considerations for next ANU running
10. Any other business
11. Date of next meeting

W.N. Catford (Secretary) 4 January 1994